Given that f(x)=2^2-11 and g(x) = 7x+1, find (fg) -(1/7), if it exists
Answer by Alan3354(69443)

Given that f(x)=2^2-11 and g(x) = 7x+1, find (fg) -(1/7), if it exists
---------
f(x) has no x term
f(x) = -7
-----
g(-1/7) = 7*(-1/7) + 1 = 0
f*g(-1/7) = 0